# On Fock covariance for product systems and the reduced Hao–Ng isomorphism problem by discrete actions

### v1

@@ 3.3. The product system of a semigroup
be a set of neutral words such that $K(\alpha_{i})=\mathbf{x}_{i}$ for $i=1,\dots,n$. Moreover, from the comments above we may pick $\mu_{i}\in\mathbb{C}$ such that $t_{\ast}(b_{\mathbf{x}_{i}})=\mu_{i}\dot{w}_{\alpha_{i}}$ for $i=1,\dots,n$. As the projections $\dot{w}_{\alpha_{i}}$ commute, by the Gel’fand-Naimark Theorem, and considering a unit decomposition (see also [34, Lemma 3.13]), there exists a subset $B$ of $F$ such that
@@
Since $b=\sum_{i=1}^{n}\mu_{i}\dot{w}_{\alpha_{i}}$, by the form of the projection $Q_{B}$ we have that
@@
where $\beta_{j}$ is a concatenation of the words $\alpha_{i}\in B$ and $\alpha_{j}$, gives that
@@
By construction, for any $j$ with $\alpha_{j}\notin B$ we have that
@@ 4.1. Operator algebras
By an operator algebra ${\mathfrak{A}}$ we will mean a norm-closed subalgebra of some ${\mathcal{B}}(H)$. Every operator algebra attains a C*-cover, i.e., a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to{\mathcal{C}}$ such that ${\mathcal{C}}=\mathrm{C}^{*}(\iota({\mathfrak{A}}))$. The C*-envelope $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ is the co-universal C*-cover of ${\mathfrak{A}}$, i.e., there exists a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that for any other C*-cover $j\colon{\mathfrak{A}}\to{\mathcal{C}}$ there exists a unique *-epimorphism $\Phi\colon{\mathcal{C}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that $\Phi\circ j=\iota$. The reader is addressed to [5, 44] for the full details.
The existence of the C*-envelope was established by Hamana [19] through the existence of the injective envelope. An independent proof was established by Dritschel and McCullough [14] through the existence of maximal dilations. Recall that a homomorphism $\phi\colon{\mathfrak{A}}\to{\mathcal{B}}(K)$ is called a dilation of $\iota\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ if $H\subseteq K$ and $\iota(\cdot)=P_{H}\phi(\cdot)|_{H}$. A dilation is called maximal if it only attains trivial dilations, i.e., dilations by orthogonal summands. As an application of the existence of maximal dilations, Dritschel and McCullough [14] show that the C*-envelope of ${\mathfrak{A}}$ is the C*-algebra generated by a maximal dilation. Arveson [3] has shown that a representation of ${\mathfrak{A}}$ is maximal if and only if it has a unique extension to a completely positive map on $\mathrm{C}^{*}({\mathfrak{A}})$ that is a *-homomorphism. The proof in [3] refers to operator systems, but it can be adapted to the operator algebras category; the reader is directed to [13] for the full details.
@@ 4.2. Strong covariant representations
We review the key elements of the strong covariant C*-algebra of a product system established by Sehnem [47, 48], and its Fell bundle considered by Dor-On, the first author, Katsoulis, Laca and Li [11].
@@
For a product system $X$ we will write ${\mathcal{T}}_{\lambda}(X)^{+}$ for the tensor algebra of $X$, i.e., for the norm-closed subalgebra generated by $\{\lambda_{p}(X_{p})\}_{p\in P}$ inside ${\mathcal{T}}_{\lambda}(X)$. We will require the following consequence of [48, Corollary 3.5]. We note that although only product systems in the sense of Fowler are considered in [48], the arguments apply in our setting as well.
@@
Let the canonical *-epimorphism $\Phi\colon{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\to\mathrm{C}^{*}(t)$. Since $t$ is equivariant, then $\Phi$ induces a *-epimorphism between the reduced C*-algebras of the induced Fell bundles. However, since the coaction on $\mathrm{C}^{*}(t)$ is normal, it follows that $\mathrm{C}^{*}(t)$ coincides with the reduced C*-algebra of its induced Fell bundle. Hence there exists a canonical *-epimorphism $\Phi_{\lambda}\colon{\mathcal{T}}_{\lambda}(X)\to\mathrm{C}^{*}(t)$, which is injective on $A$ since $t$ is injective. By [48, Corollary 3.5] it follows that the restriction of $\Phi_{\lambda}$ on ${\mathcal{T}}_{\lambda}(X)^{+}$ is completely isometric. ∎
Since $A\times_{X,\lambda}P$ satisfies the conditions of Corollary 4.1, it is a C*-cover of ${\mathcal{T}}_{\lambda}(X)^{+}$. In [48, Theorem 5.1], Sehnem establishes that $A\times_{X,\lambda}P$ is the C*-envelope of ${\mathcal{T}}_{\lambda}(X)^{+}$.
@@ 4.3. Crossed products
Let ${\mathfrak{H}}$ be a locally compact group that acts on an operator algebra ${\mathfrak{A}}$ by completely isometric automorphisms. The action of ${\mathfrak{H}}$ then extends also to a group action $\dot{\alpha}$ of the C*-envelope of ${\mathfrak{A}}$, and we can form the reduced crossed product $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$. By considering the copy of ${\mathfrak{A}}$ inside $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, the reduced crossed product ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is defined as the norm-closed subalgebra of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$ generated by the ${\mathfrak{A}}$-valued functions [27, Definition 3.17]. One of the main questions in the theory is whether this inclusion induces a canonical *-isomorphism with the C*-envelope of ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, i.e., whether it holds that
@@
We identify ${\mathfrak{A}}$ with its copy in $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$. Let $\pi\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ be a maximal completely isometric representation of ${\mathfrak{A}}$. Then $\pi$ has a unique extension to a faithful *-representation on $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, denoted by the same symbol. Therefore $\overline{\pi}\rtimes U$ is a faithful *-representation of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$, and thus its restriction on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is completely isometric. It suffices to show that $\overline{\pi}\rtimes U$ is maximal on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, as then it will follow that
@@
By using the unique extension property of $\rho$ we have that
@@
By equating the $(1,1)$-entries we get that $x=0$. On the other hand, by using the covariance in the C*-crossed product we can write
@@
Then a similar computation gives that
@@
By equating the $(1,1)$-entries we get that $y=0$. Therefore $\rho$ is a trivial dilation, as required. ∎
@@ 4.4. Hao–Ng
Let $\pi$ be a faithful *-representation of ${\mathcal{T}}_{\lambda}(X)$ in some ${\mathcal{B}}(H)$ and consider the faithful *-representation $\overline{\pi}\rtimes U$ on $\ell^{2}({\mathfrak{H}},H)$ that gives rise to ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Since $\alpha$ restricts to an action on ${\mathcal{T}}_{\lambda}(X)^{+}$, from [27, Corollary 3.16] we have a canonical completely isometric copy of the crossed product ${\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ in ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. We can now consider the family $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ of the subspaces
@@
For condition (i) we have by definition that
@@
which is a C*-algebra. The other two conditions follow by the fact that $\alpha_{{\mathfrak{h}}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, the covariant relations of the crossed product and that $\lambda$ is a product system representation. Indeed, for condition (ii) we have that
@@
By considering finite linear combinations and their norm-limits we obtain that
@@
Due to the properties of the generalised gauge action we have a canonical identification of the $\mathbf{K}_{\bullet}$-cores. Note that, since every $\lambda_{p}(X_{p})$ is $\alpha$-invariant, then so is every C*-subalgebra $\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}$ for ${\mathbf{x}}\in{\mathcal{J}}$. By construction we have that
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$ and let the identity representation $\iota^{\rtimes}\colon X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. If ${\mathbf{x}}\in{\mathcal{J}}$, then
@@
We will make the standard identification of ${\mathcal{T}}_{\lambda}(X)$ with $\overline{\pi}({\mathcal{T}}_{\lambda}(X))$, and of ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ with $\mathrm{C}^{*}(\overline{\pi},U)$. In order to make a distinction we write $e_{G}$ for the unit of $G$ and $e_{\mathfrak{H}}$ for the unit of ${\mathfrak{H}}$. For the first part note that $\iota^{\rtimes}_{\ast}(\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}})=\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}$ since $\iota^{\rtimes}$ is a representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover by checking the generators we have that
@@
By the covariance relation of $(\overline{\pi},U)$, and the fact that $\alpha_{\mathfrak{h}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, we can write
@@
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some appropriate choice of $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. Therefore we have that
@@
By considering linear combinations and their norm-limits we have that
@@
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. By definition we have that
@@
We will show that the identity representation satisfies the conditions of Theorem 3.2. By definition it is injective. Moreover, by [50, Lemma 7.16] we have that
@@
Since $b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}}\in\mathbf{K}_{{\mathbf{x}}_{i}\cap{\mathbf{x}}_{j},\hat{t}^{\rtimes}_{\ast}}$, by Proposition 4.4 we have that
@@
By using that ${\mathcal{F}}$ is $\cap$-closed we obtain that
@@
Since $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, by using the assumption we have that $c_{r}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$ for all $\xi_{r}\in X_{r}$. Thus we get
@@
where we used that ${\mathcal{T}}_{\lambda}(X)$ is in the multiplicative domain of $E_{\mathfrak{H}}$. Consequently, we get that $E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})^{1/2}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$ and in particular
@@
Since $\lambda$ satisfies condition (ii) of Theorem 3.2 we get that $\sum_{i,j=1}^{n}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})=0$, and therefore
@@
By [48, Theorem 5.1] this is equivalent to asking for a canonical *-isomorphism
@@
By Proposition 4.5 the identity representation is an equivariant Fock covariant injective representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover, by [50, Lemma 7.16] we have that
@@
the required completely isometric isomorphism is induced by Corollary 4.1. By Proposition 4.2 we then have that
@@
and [48, Theorem 5.1] finishes the proof. ∎

### v2

@@ 3.3. The product system of a semigroup
be a set of neutral words such that $K(\alpha_{i})=\mathbf{x}_{i}$ for $i=1,\dots,n$. Moreover, from the comments above we may pick $\mu_{i}\in\mathbb{C}$ such that $t_{\ast}(b_{\mathbf{x}_{i}})=\mu_{i}\dot{w}_{\alpha_{i}}$ for $i=1,\dots,n$. As the projections $\dot{w}_{\alpha_{i}}$ commute, by the Gel’fand-Naimark Theorem, and considering a unit decomposition (see also [38, Lemma 3.13]), there exists a subset $B$ of $F$ such that
@@
Since $b=\sum_{i=1}^{n}\mu_{i}\dot{w}_{\alpha_{i}}$, by the form of the projection $Q_{B}$ we have
@@
where $\beta_{j}$ is a concatenation of the words $\alpha_{i}\in B$ and $\alpha_{j}$, gives
@@
By construction, for any $j$ with $\alpha_{j}\notin B$ we have
@@ 4.1. Operator algebras
For the general theory of nonselfadjoint operator algebras and full details, the reader is addressed to [5, 48]. By an operator algebra ${\mathfrak{A}}$ we will mean a norm-closed subalgebra of some ${\mathcal{B}}(H)$. Every operator algebra attains a C*-cover, i.e., a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to{\mathcal{C}}$ such that ${\mathcal{C}}=\mathrm{C}^{*}(\iota({\mathfrak{A}}))$. The C*-envelope $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ is the co-universal C*-cover of ${\mathfrak{A}}$, i.e., there exists a completely isometric homomorphism $\iota\colon{\mathfrak{A}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that for any other C*-cover $j\colon{\mathfrak{A}}\to{\mathcal{C}}$ there exists a unique *-epimorphism $\Phi\colon{\mathcal{C}}\to\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$ such that $\Phi\circ j=\iota$.
The existence of the C*-envelope was established by Hamana [22] through the existence of the injective envelope. An independent proof was established by Dritschel and McCullough [16] through the existence of maximal dilations. Recall that a homomorphism $\phi\colon{\mathfrak{A}}\to{\mathcal{B}}(K)$ is called a dilation of $\iota\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ if $H\subseteq K$ and $\iota(\cdot)=P_{H}\phi(\cdot)|_{H}$. A dilation is called maximal if it only attains trivial dilations, i.e., dilations by orthogonal summands. Dritschel and McCullough show in [16, Theorem 1.2] that every (completely isometric) representation of an operator algebra ${\mathfrak{A}}$ admits a (resp. completely isometric) maximal dilation. Moreover in [16, Theorem 4.1] they obtain that the C*-algebra of a maximal completely isometric representation satisfies the co-universal property of the C*-envelope. In a simplified version, Arveson [3] has shown that a representation of ${\mathfrak{A}}$ is maximal if and only if it has a unique extension to a completely positive map on $\mathrm{C}^{*}({\mathfrak{A}})$ that is a *-homomorphism. The proof in [3] refers to operator systems, but it can be adapted to the operator algebras category; the reader is directed to [15, Section 2] and [51, Proposition 4.8] for the full details.
@@ 4.2. Strong covariant representations
We review the key elements of the strong covariant C*-algebra of a product system established by Sehnem [52, 53], and its Fell bundle considered by Dor-On, the first author, Katsoulis, Laca and Li [13]. The strong covariant algebra of Sehnem [52] as well as the reduced variant considered in [13] encodes the previously known constructions for a co-universal or a Cuntz-type C*-algebra, see [13, 52, 53] and the references therein for these connections.
@@
For a product system $X$ we will write ${\mathcal{T}}_{\lambda}(X)^{+}$ for the tensor algebra of $X$, i.e., for the norm-closed subalgebra generated by $\{\lambda_{p}(X_{p})\}_{p\in P}$ inside ${\mathcal{T}}_{\lambda}(X)$. We will require the following consequence of [53, Corollary 3.5]. We note that although only product systems in the sense of Fowler are considered in [53], the arguments apply in our setting as well.
@@
Let the canonical *-epimorphism $\Phi\colon{\mathcal{T}}_{\operatorname{c}}^{\operatorname{F}}(X)\to\mathrm{C}^{*}(t)$. Since $t$ is equivariant, then $\Phi$ induces a *-epimorphism between the reduced C*-algebras of the induced Fell bundles. However, since the coaction on $\mathrm{C}^{*}(t)$ is normal, it follows that $\mathrm{C}^{*}(t)$ coincides with the reduced C*-algebra of its induced Fell bundle. Hence there exists a canonical *-epimorphism $\Phi_{\lambda}\colon{\mathcal{T}}_{\lambda}(X)\to\mathrm{C}^{*}(t)$, which is injective on $A$ since $t$ is injective. By [53, Corollary 3.5] it follows that the restriction of $\Phi_{\lambda}$ on ${\mathcal{T}}_{\lambda}(X)^{+}$ is completely isometric. ∎
Since $A\times_{X,\lambda}P$ satisfies the conditions of Corollary 4.1, it is a C*-cover of ${\mathcal{T}}_{\lambda}(X)^{+}$. In [53, Theorem 5.1], Sehnem establishes that $A\times_{X,\lambda}P$ is the C*-envelope of ${\mathcal{T}}_{\lambda}(X)^{+}$.
@@ 4.3. Crossed products
Let ${\mathfrak{H}}$ be a locally compact group that acts on an operator algebra ${\mathfrak{A}}$ by completely isometric automorphisms. The action of ${\mathfrak{H}}$ then extends also to a group action $\dot{\alpha}$ of the C*-envelope of ${\mathfrak{A}}$, and we can form the reduced crossed product $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$. By considering the copy of ${\mathfrak{A}}$ inside $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, the reduced crossed product ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is defined as the norm-closed subalgebra of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$ generated by the ${\mathfrak{A}}$-valued functions [30, Definition 3.17]. One of the main questions in the theory is whether this inclusion induces a canonical *-isomorphism with the C*-envelope of ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, i.e., whether we have
@@
We identify ${\mathfrak{A}}$ with its copy in $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$. Let $\pi\colon{\mathfrak{A}}\to{\mathcal{B}}(H)$ be a maximal completely isometric representation of ${\mathfrak{A}}$. By [16] (see also the relevant comments in Subsection 4.1) the map $\pi$ has a unique extension to a faithful *-representation on $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})$, denoted by the same symbol. Therefore $\overline{\pi}\rtimes U$ is a faithful *-representation of $\mathrm{C}^{*}_{\textup{env}}({\mathfrak{A}})\rtimes_{\dot{\alpha},\lambda}{\mathfrak{H}}$, and thus its restriction on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ is completely isometric. By [16] (see also the relevant comments in Subsection 4.1), it suffices to show that $\overline{\pi}\rtimes U$ is maximal on ${\mathfrak{A}}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$, as then it will follow that
@@
By using the unique extension property of $\rho$ we have
@@
By equating the $(1,1)$-entries we get $x=0$. On the other hand, by using the covariance in the C*-crossed product we can write
@@
Then a similar computation gives
@@
By equating the $(1,1)$-entries we get $y=0$. Therefore $\rho$ is a trivial dilation, as required. ∎
@@ 4.4. Hao–Ng
Let $\pi$ be a faithful *-representation of ${\mathcal{T}}_{\lambda}(X)$ in some ${\mathcal{B}}(H)$ and consider the faithful *-representation $\overline{\pi}\rtimes U$ on $\ell^{2}({\mathfrak{H}},H)$ that gives rise to ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Since $\alpha$ restricts to an action on ${\mathcal{T}}_{\lambda}(X)^{+}$, from [30, Corollary 3.16] we have a canonical completely isometric copy of the crossed product ${\mathcal{T}}_{\lambda}(X)^{+}\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ in ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. We can now consider the family $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ of the subspaces
@@
For condition (i) by definition we have
@@
which is a C*-algebra. The other two conditions follow by the fact that we have the equality $\alpha_{{\mathfrak{h}}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, the covariant relations of the crossed product and that $\lambda$ is a product system representation. Indeed, for condition (ii) we have
@@
Due to the properties of the generalised gauge action we have a canonical identification of the $\mathbf{K}_{\bullet}$-cores. Note that, since every $\lambda_{p}(X_{p})$ is $\alpha$-invariant, then so is every C*-subalgebra $\mathbf{K}_{{\mathbf{x}},\lambda_{\ast}}$ for ${\mathbf{x}}\in{\mathcal{J}}$. By construction we have
@@
Let $P$ be a unital subsemigroup of a discrete group $G$ and let $X$ be a product system over $P$. Let $\alpha$ be a generalised gauge action of a discrete group ${\mathfrak{H}}$ on ${\mathcal{T}}_{\lambda}(X)$ and let $\iota^{\rtimes}\colon X\rtimes_{\alpha,\lambda}{\mathfrak{H}}\to{\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ be the identity representation. If ${\mathbf{x}}\in{\mathcal{J}}$, then
@@
We will make the standard identification of ${\mathcal{T}}_{\lambda}(X)$ with $\overline{\pi}({\mathcal{T}}_{\lambda}(X))$, and of ${\mathcal{T}}_{\lambda}(X)\rtimes_{\alpha,\lambda}{\mathfrak{H}}$ with $\mathrm{C}^{*}(\overline{\pi},U)$. In order to make a distinction we write $e_{G}$ for the unit of $G$ and $e_{\mathfrak{H}}$ for the unit of ${\mathfrak{H}}$. For the first part note that we have $\iota^{\rtimes}_{\ast}(\mathbf{K}_{{\mathbf{x}},\hat{t}^{\rtimes}_{\ast}})=\mathbf{K}_{{\mathbf{x}},\iota^{\rtimes}_{\ast}}$ since $\iota^{\rtimes}$ is a representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover by checking the generators we obtain
@@
By the covariance relation of $(\overline{\pi},U)$, and the fact that we have the equality $\alpha_{\mathfrak{h}}(\lambda_{p}(X_{p}))=\lambda_{p}(X_{p})$ for all $p\in P$ and ${\mathfrak{h}}\in{\mathfrak{H}}$, we can write
@@
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some appropriate choice of $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. Therefore we have
@@
By considering linear combinations and their norm-limits we have
@@
for ${\mathfrak{h}}:={\mathfrak{h}}_{1}^{-\varepsilon}{\mathfrak{h}}_{2}\cdots{\mathfrak{h}}_{2n-1}^{-1}{\mathfrak{h}}_{2n}^{\varepsilon^{\prime}}$, and some $\eta_{p_{i}}\in X_{p_{i}}$ and $\eta_{q_{i}}\in X_{q_{i}}$ for $i=1,\dots,n$. By definition we have
@@
We will show that the identity representation satisfies the conditions of Theorem 3.2. By definition it is injective. Moreover, by [55, Lemma 7.16] we have
@@
Since $b_{\mathbf{x}_{i}}^{*}b_{\mathbf{x}_{j}}\in\mathbf{K}_{{\mathbf{x}}_{i}\cap{\mathbf{x}}_{j},\hat{t}^{\rtimes}_{\ast}}$, by Proposition 4.4 we have
@@
By using that ${\mathcal{F}}$ is $\cap$-closed we obtain
@@
Since $r\in\bigcup_{i=1}^{n}\mathbf{x}_{i}$, by using the assumption we have $c_{r}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$ for all $\xi_{r}\in X_{r}$. Thus we get
@@
where we used that ${\mathcal{T}}_{\lambda}(X)$ is in the multiplicative domain of $E_{\mathfrak{H}}$. Consequently, we obtain
$E_{{\mathfrak{H}}}(c_{r}^{*}c_{r})^{1/2}\overline{\pi}(\lambda_{r}(\xi_{r}))=0$
and in particular
@@
Since $\lambda$ satisfies condition (ii) of Theorem 3.2 we get $\sum_{i,j=1}^{n}\lambda_{\ast}(d_{\mathbf{x}_{i}\cap\mathbf{x}_{j}})=0$, and therefore
@@
By [53, Theorem 5.1] this is equivalent to asking for a canonical *-isomorphism
@@
By Proposition 4.5 the identity representation is an equivariant Fock covariant injective representation of $X\rtimes_{\alpha,\lambda}{\mathfrak{H}}$. Moreover, by [55, Lemma 7.16] we have
@@
the required completely isometric isomorphism is induced by Corollary 4.1. By Proposition 4.2 we then have
@@
and [53, Theorem 5.1] finishes the proof. ∎